OpenStudy (anonymous):

Medal for Matrix help. How would I solve for x? [-8 -14 -2x] = [2 -74 -6]

OpenStudy (usukidoll):

what row by column matrix is this?

OpenStudy (anonymous):

1 x 3 for both

OpenStudy (usukidoll):

so is A - [-8 -14 2x] and B =[2 -74 -6]?

OpenStudy (usukidoll):

it's like find an x that will make both matrices equal...

OpenStudy (anonymous):

Yeah, that's the issue. I have no idea how to find x in this situation. It's been a while since I've done a problem like this.

OpenStudy (usukidoll):

oh how about if we have something like this -2x=-6 then we just solve for x. Because what I am assuming is that the last entry in A and B are supposed to be equal.

OpenStudy (usukidoll):

I have done these, but with 2x2 or above

OpenStudy (anonymous):

I could give you the possible answers if that would help?

OpenStudy (anonymous):

a. x = -5.7 b. x = -17.7 c. x = 7 d. x = 6

OpenStudy (usukidoll):

yes please

OpenStudy (usukidoll):

what the O_O!

OpenStudy (anonymous):

Yeah...my reaction exactly.

OpenStudy (usukidoll):

this is so weird.. I'm calling wolfram lol

OpenStudy (anonymous):

lol okay. That's a good idea. This was the original problem... \[[-8 -14-2x] - ([-19 +17 -1] + [14 +13 -2]) = [-3 -44 -9]\]

OpenStudy (usukidoll):

T_T wait a sec.. I might have ... did you do the operation in paraenthesis first?

OpenStudy (anonymous):

Oops! Ignore the plus signs. I dunno why I put them inside the matrices. and yes I did. Was I not supposed to?

OpenStudy (usukidoll):

-.- ok hold on I think I might have see what is going.. yes you should do the ones in paraenthesis first ..

OpenStudy (loser66):

Question: how A = B??

OpenStudy (usukidoll):

that's what I was thinking too.

OpenStudy (usukidoll):

what makes A=B ? but when I was to do the problem.. if I let x = 3 that last part will be equal but the first 2 entries aren't equal

OpenStudy (loser66):

if A = B ,then -8 =2?? -14 =-74 -2x = -6 but how??

OpenStudy (usukidoll):

something is wrong with the arithmetic.

OpenStudy (usukidoll):

let's do it over... and no shortcut cheats.. now this is the original problem right? \[[-8 -14-2x] - ([-19 +17 -1] + [14 +13 -2]) = [-3 -44 -9]\]

OpenStudy (anonymous):

That wouldn't surprise me. I've been doing math since about 10 am. and yes it is.

OpenStudy (usukidoll):

\[[-8, -14,-2x] - ([-19 , 17 , -1] + [14, 13 ,-2]) = [-3, -44, -9]\]

OpenStudy (usukidoll):

let me put on my cglasses

OpenStudy (usukidoll):

now let's do the ones in the () with adding the matrices.

OpenStudy (usukidoll):

\[- ([-19 , 17 , -1] + [14, 13 ,-2]) \]

OpenStudy (usukidoll):

\[- ([-19+14 , 17+13 , -1-2] )\]\[-[-5,30,-3] \rightarrow [5,-30,3]\]

OpenStudy (usukidoll):

\[[-8 -14-2x] - [5, -30, 3] \rightarrow [-8-5-14-(-30),-2x-3]\] are you sure it's not [8, -14, 2x] ?

OpenStudy (usukidoll):

\[[8 -14-2x] - [5, -30, 3] \rightarrow [8-5-14-(-30),-2x-3] \rightarrow [3, 14+30,-2x-3] \rightarrow [3, 44, -2x-3\] \[ \rightarrow [3, 14+30,-2x-3] \rightarrow [3, 44, -2x-3]\]

OpenStudy (usukidoll):

yay I think I got what x is !

OpenStudy (usukidoll):

on the last entry we have to solve for x so now we have \[ [3, 44, -2x-3] = [3, 44, 9]\]

OpenStudy (anonymous):

Really? (sorry, I'm doing this an something else at the same time)

OpenStudy (usukidoll):

wait a sec... I think there are typos

OpenStudy (usukidoll):

I think.. let me do this on paper.. there is a typo somewhere on your end or something.

OpenStudy (anonymous):

Okay.

OpenStudy (usukidoll):

-_- I got x = 3. who are these people giving these choices anyway?

OpenStudy (usukidoll):

\[[-8, -14, 2x] +[ 5 -30 ,3] = [-8+5, -14-30, 2x+3] = [-3, -44, 2x+3]\]

OpenStudy (anonymous):

It's just random questions I'm given for practice. At the moment, I am unable to view the question, but as soon as I can, I'll let you know if there were any typos in my copy.

OpenStudy (usukidoll):

\[[-3, -44, 2x+3] = [-3, -44, -9]\] there could be typos on my end but I got it... ok THIS TIME I AM CERTAIN I GOT IT!

OpenStudy (usukidoll):

we just have to focus on the last equation because our first two entries are equal so we now solve for x for that last entry 2x+3=-9 2x =-9-3 2x=-12 x =- 6 there's a typo in your multiple choice

OpenStudy (usukidoll):

yup x = -6 \[[-3, -44, 2(-6)+3] = [-3, -44, -12+3] = [-3, -44, -9]\]

OpenStudy (usukidoll):

yay finally it's over. can I get a medal for this 1 x 3 horror festival? I'm not use to adding and solving like that XD

OpenStudy (usukidoll):

-2x+3 = -9 doh I dropped a sign. It's D -2x = -9-3 -2x=-12 x=6

OpenStudy (anonymous):

You're awesome! :D Hopefully I'll never get another question like this again, but thank you so much for all your help.

OpenStudy (usukidoll):

that question sucked... I was never given like this and I'm used to adding vertically!

OpenStudy (anonymous):

Honestly, I've never had one like this either. The moment I looked at it, my brain went blank.

OpenStudy (usukidoll):

it was tedious. doable but tedious... I think the latex threw me off too. That's why I got so many answers and what not.

OpenStudy (usukidoll):

Cleaned up version \[[-8 -14-2x] - ([-19 +17 -1] + [14 +13 -2]) = [-3 -44 -9]\] \[\[[-8 -14-2x] - ([-19+14 , 17+13 , -1-2] ) = [-3 -44 -9]\] \[\[[-8 -14-2x] - ([-5, 30 , -3] ) = [-3 -44 -9]\] \[\[[-8 -14-2x] +([5, -30 , 3] ) = [-3 -44 -9]\] \[\[[-8+5, -14-30,-2x+3] ) = [-3 -44 -9]\] \[\[[-3, -44,-2x+3] ) = [-3 -44 -9]\] \[-2x+3=-9 \] \[-2x+3-3=-9-3 \] \[-2x=-12 \] \[\frac{-2x}{-2}=\frac{-12}{-2}\] \[x=6\] \[\[[-3, -44,-2(6)+3] ) = [-3 -44 -9]\] \[\[[-3, -44,-12+3] ) = [-3 -44 -9]\] \[\[[-3, -44,-9] ) = [-3 -44 -9]\]